# Choose Your Battles: Distributed Learning Over Multiple Tug of War Games

### v1

@@ 3 Tug-of-Peace Algorithms
We then present the Meta-ToP algorithm for the setting of Meta-ToW games. We borrow intuition from the single game setting and use the $1$-bit communication for players to switch between game configurations till they find a configuration under which the QoS requirements are feasible. We show that in a finite number of game switches, our algorithm converges to a game configuration and action profile that satisfies with QoS requirement with probability 1.
@@ 3.1 Tug-of-Peace Algorithm
Now, let us discuss the iteration performed by each player. For simplicity, first consider a case where there is no noise, so each player observes their exact reward, i.e., $y_{n}(t)=u_{n}(\boldsymbol{x}(t))$. Then the idea behind the algorithm is that each player $n$ starts with action 0 and as they receive a reward lower than their QoS requirement, they increase their action. This increase is at a rate proportional to its ‘dissatisfaction’, i.e., how far their reward is from the QoS requirement. By definition of ToW games, this causes a drop in reward received by other players which encourages them to increase their actions. This ‘cooperative’ increase in actions eventually leads to convergence to an equilibrium point.
Even if there exists a feasible point in the interior, the noise can cause a player to reach the boundary and get stuck there. To avoid this, when a player tries to exceed the boundary, it sends a signal to other players stating that it might be stuck at the boundary. On receiving such a signal, all players return back to the action 0. This resets the iteration and brings it to a point that is in the domain of attraction of an equilibrium where the QoS condition is satisfied. With a lower starting stepsize after the reset, as required by Assumption 3, the probability of reaching the boundary due to noise is lower.
The signal that indicates a player is stuck at the boundary carries $1$-bit of information. We show that these resets happen only a finite number of times with probability 1. Hence, the communication overhead of Algorithm 1 is negligible. Implementing this communication is application-dependent. For example, communication between devices is more obvious in sensor networks than in power control and task allocation.
We need to choose appropriate stepsizes to deal with noise and the resets, specified in the following standard assumption:
@@ 3.2 Fully Distributed Tug-of-Peace Algorithm
Under assumptions 1, 2, and 3, the iterates of Algorithm 2 converge with probability 1 to a point. The iterates of Algorithm 2 converge to $\boldsymbol{x}_{*}$, as defined in Definition 3, with probability $1-\varepsilon(\{\eta\})$ where $\varepsilon(\{\eta\})\rightarrow 0$ as $\eta(0)\rightarrow 0$.
@@ 3.3 Meta Tug-of-Peace Algorithm
We now present the Meta Tug-of-Peace (Meta-ToP) algorithm, a distributed algorithm that ensures convergence to an action profile where all players achieve their QoS requirements in Meta-ToW games. Similar to the ToP algorithm, this algorithm has the two key concepts of ‘cooperative’ increase in actions and boundary signals. But here the boundary signals play two roles: avoiding getting stuck at the boundary, similar to ToP, and switching game configurations. If a boundary signal is broadcast, it is a possible indicator that the QoS guarantees are not feasible under the current game configuration, and hence the players change the configuration. It is always possible that the configuration was feasible, and a player reached the boundary due to noise. But our algorithm ensures that each configuration will be tried infinitely often up until the QoS requirements are met.
Meta-ToP works as follows: suppose player $n$ reaches the boundary $B_{n}$ at time $t$. It broadcasts $s=1$ to all players in the same game, i.e., all players $m$ with $g_{m}(t)=g_{n}(t)$ and signals $r=1$ to all players in all games. Each player that received signal $s=1$ switches a game with probability $\rho>0$. If a player $m$ decides to switch a game, they choose a game uniformly at random from $\mathcal{G}\setminus\{g_{m}(t)\}$. Players which only receive signal $r=1$ perform the same procedure but with probability $\varphi>0$ instead of $\rho$. We call this procedure the game switch mechanism. Each player $m$, irrespective of whether they switched their game or not, resets their action to $x_{m}(t+1)=0$ when they receive signal $r=1$.
@@
Our following result gives the convergence guarantees for the MC-ToP algorithm:
@@ 4.1 Analysis of the ToP Algorithm for a Single ToW Game
By definition of a ToW game, this ODE satisfies the property $\frac{\partial h_{n}}{x_{m}}>0$ for all $n\neq m\in\mathcal{N}$. Such an ODE is called a cooperative ODE [22, 23]. We have already assumed that there exists an equilibrium point for this ODE in the region $\mathcal{X}^{o}$ (Assumption 2(a)). Then this class of ODEs have certain desirable convergence properties, which we restate in the following lemma:
@@
    [24, Theorem 2.1] For initial conditions in an open dense set, the solutions of (4) converge to an equilibrium.
@@
    [12, Theorem 5.6] There exists a minimal equilibrium $\boldsymbol{x}_{*}$ of (4) such that any other equilibria $\hat{\boldsymbol{x}}$ satisfies $x_{*_{n}}\leq\hat{x}_{n}$ for all $n\in\mathcal{N}$.
@@
The equilibrium point $\boldsymbol{x}_{*}$ is a stable equilibrium point for the ODE (4) with probability 1. Moreover, it is a Locally Asymptotically Stable Equilibrium (LASE) [11, B.3].
@@
Let $u(\cdot)$ be the vector function $(u_{1}(\cdot),\ldots,u_{N}(\cdot))^{T}$. Then $Dh(\boldsymbol{x})=-Du(\boldsymbol{x})$ for all $\boldsymbol{x}$, where $Dh(\boldsymbol{x})$ and $Du(\boldsymbol{x})$ denote the Jacobians of functions $h$ and $u$, respectively, computed at $\boldsymbol{x}$. Then Sard’s Lemma [25] states that the image of points for which the Jacobian $Du(\cdot)$ is nonsingular has Lebesgue measure zero, i.e., the set of points $\mathbf{z}$ such that $u(\boldsymbol{x})=\mathbf{z}$ and $Du(\boldsymbol{x})=0$ has Lebesgue measure zero. Let this set with Lebesgue measure zero be denoted by $\mathcal{Z}$. Since each $\boldsymbol{\overline{\lambda}}$ is chosen uniformly from the set $\lambda_{n},\lambda_{n}+\delta]$, $\boldsymbol{\overline{\lambda}}\notin\mathcal{Z}$ with probability 1. This implies that $Dh(\boldsymbol{x}_{*})$ is nonsingular a.s. which implies that the equilibrium point $\boldsymbol{x}_{*}$ is a.s. isolated. Using Assumption 2(b), the Jacobian $Dh(\boldsymbol{x}_{*})$ does not have purely imaginary eigenvalues a.s. and hence the equilibrium $\boldsymbol{x}_{*}$ is hyperbolic a.s. Theorem 4.1.1 from [26] states that if the minimal equilibrium $\boldsymbol{x}_{*}$ is hyperbolic, then it is a stable equilibrium. We know that the solutions of (4) converge, and hence $\boldsymbol{x}_{*}$ is also a LASE. ∎
Lemma 5 implies that any solution of the ODE initiated in $0,\boldsymbol{x}_{*}]$ will converge to the equilibrium point $\boldsymbol{x}_{*}$ and hence $\boldsymbol{x}=\mathbf{0}_{N}$ is in the domain of attraction of the LASE $\boldsymbol{x}_{*}$. We use this fact later in Lemma 7. The next lemma, based on [27, Section 5.1], argues about the convergence of the iteration (3) in the FDToP algorithm to an ODE and therefore to equilibria.
@@
    with probability 1. Here $b(\boldsymbol{x}(t))=\textbf{0}$ in $0,B_{1})\times\ldots\times[0,B_{N})$ and $b_{n}(\boldsymbol{x}(t))=-h_{n}(\boldsymbol{x}(t))$ if $x_{n}(t)=B_{n}$ and $h_{n}(\boldsymbol{x}(t))>0$.
@@
For a system satisfying assumptions 2-3, and large enough $t^{\prime},T$, and small enough $\epsilon$, the iterates of Algorithm 1 and 2 satisfy the following:
@@
Finally, through an extension of Borel-Cantelli Lemma [29, Corollary 5.29], we have that with probability 1:
@@
Let $\tau$ denote the last such reset. Then for $t>\tau$, the iterates of (2) always stay in the interior of $\mathcal{X}$. Then the iterates a.s. asymptotically track the solutions of the ODE (4). Hence actions of all players converge to an equilibrium point of the form $\hat{\boldsymbol{x}}$ which satisfies $u_{n}(\hat{x}_{n})=\overline{\lambda}_{n}$ for all players $n$. ∎
Proof of Theorem 1 (b).
The algorithm is initialized at $\boldsymbol{x}(0)=\mathbf{0}_{N}$, so for a small enough $\eta(0)$, Lemma 7 can be applied for $t^{\prime}=0$. Thus, with probability $1-\phi(0)$, the iterates stay in the $\epsilon$-vicinity of $\boldsymbol{x}_{*}$ from some $T$ onwards and converge to $\boldsymbol{x}_{*}$ as $\boldsymbol{x}_{*}$ is a LASE. Note that $\phi(0)\rightarrow 0$ as $\eta(0)\rightarrow 0$. ∎
@@
Lemma 6 (b) implies that the iterates of Algorithm 2 converge to a point. As $\boldsymbol{x}(0)=\mathbf{0}_{N}$, for small enough $\eta(0)$, Lemma 7 can be applied for $t^{\prime}=0$. And hence with probability at least $1-\phi(0)$, the iterates stay in the $\epsilon$-vicinity of $\boldsymbol{x}_{*}$ from some $T$ onwards and converge to $\boldsymbol{x}_{*}$ as $\boldsymbol{x}_{*}$ is a LASE. Note that $\phi(0)\rightarrow 0$ as $\eta(0)\rightarrow 0$. ∎
@@ 4.2 Analysis of the Meta-ToP Algorithm
Finally, through an extension of Borel-Cantelli Lemma [29, Corollary 5.29], we have that with probability 1:
@@
5 Applications
We now study three problems that can be modeled using Meta-ToW games and the QoS achievability problem.
5.1 Multi-Channel Power Control in Wireless Networks
In power control, the players are $N$ transmitter-receiver pairs, so transmitter $n$ wishes to transmit to receiver $n$. There are $K$ channels and at each timestep, each player $n$ decides the channel $g_{n}$ to transmit over. For each player $n$, the action $x_{n}$ denotes the transmission power of player $n$ in channel $g_{n}$. The interference experienced by receiver $n$ is given by
$I_{n}(\boldsymbol{g},\boldsymbol{x})=\sum_{m\neq n:g_{m}=g_{n}}c_{m,n}x_{m},$
$\frac{\partial u_{n}}{\partial x_{m}}=-\frac{c_{n,n}c_{m,n}x_{n}}{(N_{0}+I_{n}(\boldsymbol{g},\boldsymbol{x}))^{2}}<0$
for all $m\neq n$ such that $g_{m}=g_{n}$ and $\boldsymbol{x}\in\mathcal{X}^{o}$. There are multiple different assumptions on $\boldsymbol{\lambda}$ and the matrix $C=[c_{m,n}]$ in the literature [2, 1] which all result in the existence of a point that satisfies the QoS requirement. Taking the boundary $B_{n}$ to be large enough for each player $n$ would ensure that the QoS is achieved in the bounded region $0,B_{n}]$. Moreover, in case of a single channel (or equivalently a single game in our framework) these assumptions also result in the uniqueness of equilibrium point $\hat{\boldsymbol{x}}$ which satisfies $u_{n}(\boldsymbol{x})=\lambda_{n}$ for all players $n$. Trivially this point is also the minimal equilibrium point.
In wireless networks, the locations of the devices, and thus the channel gains, are typically modeled as random [1]. Then, there is no need for Assumption 2 as the randomness in the channel gain matrix $C$ has a similar effect on its eigenvalues.
5.2 Distributed Task Allocation with Cooperative Agents
Consider $N$ agents who collaborate to perform $K$ tasks [31, 32]. Each player decides which task to work on and how much effort they make on this task. The total utility obtained from a task depends on all the agent’s efforts working on that task, which can be thought of as the value of the progress rate of this task. This total utility is then divided between the agents working on that task, which represents their marginal contribution to the total value attributed to the progress in the task. This total utility is monotonically increasing and concave with respect to the total effort spent on that task. Progress rate typically increases sublinearly in effort or linearly at best if there is no overhead in collaboration. Then, the value of increased progress usually has diminishing returns. For example, computing a job in one week instead of two is a much more significant boost than computing a job in one minute instead of two. This reward structure incentivizes agents to choose tasks where their marginal contribution is more significant. This reward, which measures the local progress of an agent, can be naturally measured locally, as opposed to the global progress. Examples for applications are LLM agents, cloud computing, and multirobot manufacturing or construction.

### v2

@@ 4 Tug-of-Peace Algorithms
We then present the Meta-ToP algorithm for the setting of Meta-ToW games. We borrow intuition from the single game setting and use the $1$-bit communication for players to switch between game configurations till they find a configuration under which the QoS requirements are feasible. We show that in a finite number of game switches, our algorithm converges to a game configuration and action profile that satisfy the QoS requirement with probability 1.
@@ 4.1 Tug-of-Peace Algorithm
Now, let us discuss the iteration performed by each player. For simplicity, first consider a case where there is no noise, so each player observes their exact reward, i.e., $y_{n}(t)=u_{n}(\boldsymbol{x}(t))$. The idea behind the algorithm is that each player $n$ starts with action 0 and, upon receiving a reward below their QoS requirement, increases their action. This increase is at a rate proportional to its ‘dissatisfaction’, i.e., how far its reward is from the QoS requirement. By definition of ToW games, this causes a drop in the rewards received by other players, which encourages them to increase their actions. This ‘cooperative’ increase in actions eventually leads to convergence to an equilibrium point.
Even if there exists a feasible point in the interior, the noise can cause a player to reach the boundary and get stuck there. To avoid this, when a player tries to exceed the boundary, it sends a signal to other players stating that it might be stuck at the boundary. On receiving such a signal, all players return to the action 0. This resets the iteration and brings it to a point that is in the domain of attraction of an equilibrium where the QoS condition is satisfied. With a lower starting stepsize after the reset, as required by Assumption 3, the probability of reaching the boundary due to noise is lower.
The signal that indicates a player is stuck at the boundary carries $1$-bit of information. We show that these resets happen only a finite number of times with probability 1. Hence, the communication overhead of Algorithm 1 is negligible. Implementing this communication is application-dependent. For example, communication between devices is more obvious in sensor networks than in power control and task allocation. If agents can only talk with their neighbors on a graph, then there will be a delay of some turns until all players receive these signals, but nothing else changes in our analysis or results.
We need to choose appropriate stepsizes to deal with the noise and the resets, specified in the following standard assumption:
@@
  3. (c)
    The iterates of Algorithm 1 stay within an $\epsilon$-ball of $\boldsymbol{x}_{*}$ after $T=\mathcal{O}(\log(1/\epsilon))$ with probability $1-\varepsilon(\{\eta\})$.
@@ 4.2 Fully Distributed Tug-of-Peace Algorithm
Under assumptions 1, 2, and 3, the iterates of Algorithm 2 converge with probability 1 to a point. The iterates of Algorithm 2 converge to $\boldsymbol{x}_{*}$, as defined in Definition 3, with probability $1-\varepsilon(\{\eta\})$ where $\varepsilon(\{\eta\})\rightarrow 0$ as $\eta(0)\rightarrow 0$. Moreover, the iterates of Algorithm 2 stay within an $\epsilon$-ball of $\boldsymbol{x}_{*}$ after $T=\mathcal{O}(\log(1/\epsilon))$ with probability $1-\varepsilon(\{\eta\})$.
@@ 4.3 Meta Tug-of-Peace Algorithm
We now present the Meta Tug-of-Peace (Meta-ToP) algorithm, a distributed algorithm that ensures convergence to an action profile where all players achieve their QoS requirements in Meta-ToW games. Similar to the ToP algorithm, this algorithm has the two key concepts of ‘cooperative’ increase in actions and boundary signals. But here the boundary signals play two roles: avoiding getting stuck at the boundary, similar to ToP, and switching game configurations. If a boundary signal is broadcast, it is a possible indicator that the QoS guarantees are not feasible under the current game configuration, and hence the players change the configuration. It is always possible that the configuration was feasible, and a player reached the boundary due to noise. But our algorithm ensures that each configuration will be tried infinitely often until the QoS requirements are met.
Meta-ToP works as follows: suppose player $n$ reaches the boundary $B_{n}$ at time $t$. It broadcasts $s=1$ to all players in the same game, i.e., all players $m$ with $g_{m}(t)=g_{n}(t)$ and signals $r=1$ to all players in all games. Each player who received signal $s=1$ switches a game with probability $\rho>0$. If a player $m$ decides to switch a game, they choose a game uniformly at random from $\mathcal{G}\setminus\{g_{m}(t)\}$. Players which only receive signal $r=1$ perform the same procedure but with probability $\varphi>0$ instead of $\rho$. We call this procedure the game switch mechanism. Each player $m$, irrespective of whether they switched their game or not, resets their action to $x_{m}(t+1)=0$ when they receive signal $r=1$.
@@
Our following result gives the convergence guarantees for the Meta-ToP algorithm:
@@
Technically, any probabilities $0<\rho,\varphi<1$ guarantee that all game configurations are tried ‘sufficiently often’. Thus, the asymptotic convergence guaranteed by Theorem 3 holds for any $0<\rho,\varphi<1$. However, $\rho,\varphi$ still affect the convergence rate. Conveniently, our simulations in Section 6 show that convergence time is only mildly sensitive to these parameters.
@@ 5.1 Analysis of the ToP Algorithm for a Single ToW Game
By definition of a ToW game, this ODE satisfies $\frac{\partial h_{n}}{\partial x_{m}}>0$ for all $n\neq m\in\mathcal{N}$. Such an ODE is called a cooperative ODE [33, 23]. We have already assumed that there exists an equilibrium point for this ODE in the region $\mathcal{X}^{o}$ (Assumption 2(a)). Then this class of ODEs has certain desirable convergence properties, which we restate in the following lemma:
@@
    [22, Theorem 2.1] For initial conditions in an open dense set, the solutions of (7) converge to an equilibrium.
@@
    [21, Theorem 5.6] There exists a minimal equilibrium $\boldsymbol{x}_{*}$ of (7) such that any other equilibria $\hat{\boldsymbol{x}}$ satisfies $x_{*_{n}}\leq\hat{x}_{n}$ for all $n\in\mathcal{N}$.
@@
The equilibrium point $\boldsymbol{x}_{*}$ is a stable equilibrium point for the ODE (7) with probability 1. Moreover, it is a Locally Asymptotically Stable Equilibrium (LASE) [9, B.3].
@@
Let $u(\cdot)$ be the vector function $(u_{1}(\cdot),\ldots,u_{N}(\cdot))^{T}$. Then $Dh(\boldsymbol{x})=-Du(\boldsymbol{x})$ for all $\boldsymbol{x}$, where $Dh(\boldsymbol{x})$ and $Du(\boldsymbol{x})$ are the Jacobians of functions $h$ and $u$, respectively, evaluated at $\boldsymbol{x}$. Sard’s Lemma [32] states that the image of the set of points for which the Jacobian $Du(\cdot)$ is singular has Lebesgue measure zero, i.e., the set
$\mathcal{Z}=\{\mathbf{z}\in\mathbb{R}^{N}:\exists\boldsymbol{x}\in\mathbb{R}^{N}\;\text{s.t.}\;u(\boldsymbol{x})=\mathbf{z},\;\det(Du(\boldsymbol{x}))=0\}$
has Lebesgue measure zero. Since each coordinate of $\boldsymbol{\overline{\lambda}}$ is chosen uniformly from the set $\lambda_{n},\lambda_{n}+\delta]$, $\boldsymbol{\overline{\lambda}}\notin\mathcal{Z}$ a.s. This implies that $Dh(\boldsymbol{x}_{*})$ is nonsingular a.s. which implies that the equilibrium point $\boldsymbol{x}_{*}$ is a.s. isolated. Using Assumption 2(b), the Jacobian $Dh(\boldsymbol{x}_{*})$ does not have purely imaginary eigenvalues a.s. and hence the equilibrium $\boldsymbol{x}_{*}$ is hyperbolic a.s. Theorem 4.1.1 from [10] states that if the minimal equilibrium $\boldsymbol{x}_{*}$ is hyperbolic, then it is a stable equilibrium.
Now, since $\boldsymbol{x}_{*}$ is isolated, there exists an open neighborhood $U\in\mathcal{X}^{o}$ of $\boldsymbol{x}_{*}$ such that $\boldsymbol{x}_{*}$ is the only equilibrium in $U$. Then, because $\boldsymbol{x}_{*}$ is stable, there exists a neighborhood $V\subseteq U$ such that every solution of (7) initiated in $V$ remains in $U$ for all $t\geq 0$. Since every solution of (7) converges to an equilibrium (Lemma 1(a)), each solution starting in $V$ converges to an equilibrium in $U$. By uniqueness of the equilibrium in $U$, this limit must be $\boldsymbol{x}_{*}$. Therefore, every solution starting in $V$ converges to $\boldsymbol{x}_{*}$. Hence $\boldsymbol{x}_{*}$ is a LASE. ∎
Lemma 2 implies that any solution of the ODE initiated in $0,\boldsymbol{x}_{*}]$ will converge to the equilibrium point $\boldsymbol{x}_{*}$ and hence $\boldsymbol{x}=\mathbf{0}_{N}$ is in the domain of attraction of the LASE $\boldsymbol{x}_{*}$. We use this fact later in Lemma 4. The next lemma, based on [25, Section 5.1], argues about the convergence of the iteration (6) in the FDToP algorithm to an ODE and therefore to equilibria.
@@
    with probability 1. Here $b(\boldsymbol{x}(t))=\textbf{0}$ in $0,B_{1})\times\ldots\times[0,B_{N})$ and $b_{n}(\boldsymbol{x}(t))=-h_{n}(\boldsymbol{x}(t))$ if $x_{n}(t)=B_{n}$ and $h_{n}(\boldsymbol{x}(t))>0$ for some $n$.
@@
For a system satisfying assumptions 2-3, large enough $t^{\prime}$, and small enough $\epsilon$, there exists $T=\mathcal{O}(\log(1/\epsilon))$ such that the iterates of Algo. 1 and 2 satisfy the following:
@@
Finally, through an extension of Borel-Cantelli Lemma [12, Corollary 5.29], we have that with probability 1:
@@
Let $\tau$ denote the last such reset. Then for $t>\tau$, the iterates of (5) always stay in the interior of $\mathcal{X}$. Then the iterates a.s. asymptotically track the solutions of the ODE (7). Hence, the actions of all players converge to an equilibrium point of the form $\hat{\boldsymbol{x}}$ which satisfies $u_{n}(\hat{\boldsymbol{x}})=\overline{\lambda}_{n}$ for all players $n$. ∎
Proof of Theorem 1 (b) and (c).
The algorithm is initialized at $\boldsymbol{x}(0)=\mathbf{0}_{N}$, so for a small enough $\eta(0)$, Lemma 4 can be applied for $t^{\prime}=0$. Thus, with probability $1-\phi(0)$, the iterates stay in the $\epsilon$-vicinity of $\boldsymbol{x}_{*}$ from some $T$ onwards and converge to $\boldsymbol{x}_{*}$ as $\boldsymbol{x}_{*}$ is a LASE, where $T=\mathcal{O}(\log(1/\epsilon))$. We define $\varepsilon(\{\eta\})=\phi(0)$ and note that $\phi(0)\rightarrow 0$ as $\eta(0)\rightarrow 0$. ∎
@@
Lemma 3 (b) implies that the iterates of Algorithm 2 converge to a point. As $\boldsymbol{x}(0)=\mathbf{0}_{N}$, for small enough $\eta(0)$, Lemma 4 can be applied for $t^{\prime}=0$. And hence with probability at least $1-\phi(0)$, the iterates stay in the $\epsilon$-vicinity of $\boldsymbol{x}_{*}$ from some $T$ onwards and converge to $\boldsymbol{x}_{*}$ as $\boldsymbol{x}_{*}$ is a LASE, where $T=\mathcal{O}(\log(1/\epsilon))$. We define $\varepsilon(\{\eta\})=\phi(0)$ and note that $\phi(0)\rightarrow 0$ as $\eta(0)\rightarrow 0$. ∎
@@ 5.2 Analysis of the Meta-ToP Algorithm
Finally, through an extension of Borel-Cantelli Lemma [12, Corollary 5.29], we have that with probability 1:
@@
6 Simulations
We simulated the games from Section 3. Except for single-run plots, utility results are averaged over 100 independent runs, with the shaded region indicating the interquartile range (region between the first and third quartiles). Plots showing the percentage of converged runs depict the fraction of runs that have converged by time $t$, computed over 500 runs.